# ContraGen: A Multi-Agent Generation Framework for Enterprise Contradictions Detection

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) integrates LLMs with external sources, offering advanced capabilities for information access and decision-making. However, contradictions in retrieved evidence can result in inconsistent or untrustworthy outputs, which is especially problematic in enterprise settings where compliance, governance, and accountability are critical. Existing benchmarks for contradiction detection are limited to sentence-level analysis and do not capture the complexity of enterprise documents such as contracts, financial filings, compliance reports, or policy manuals. To address this limitation, we propose ContraGen, a contradiction-aware benchmark framework tailored to enterprise domain. The framework generates synthetic enterprise-style documents with embedded contradictions, enabling systematic evaluation of both intra-document and cross-document consistency. Automated contradiction mining is combined with human-in-the-loop validation to ensure high accuracy. Our contributions include generating realistic enterprise documents, modeling a taxonomy of contradiction types common in business processes, enabling controlled creation of self- and pairwise contradictions, developing a contradiction-aware retrieval evaluation pipeline and embedding human oversight to reflect domain-specific judgment complexity. This work establishes a foundation for more trustworthy and accountable RAG systems in enterprise information-seeking applications, where detecting and resolving contradictions is essential for reducing risk and ensuring compliance.
I Introduction
@@
Large language models (LLMs) are increasingly used across enterprises to draft policies, summarize compliance reports, generate HR documents, and support decision-making. These tasks differ from open-domain generation or short-form question answering. Enterprise text is dense, context-dependent, and often contains contradictions [^1] Contradictions may correspond to conflicting business rules, overlapping compliance guidelines, or inconsistent reporting., either introduced inadvertently or arising from evolving regulations, competing departmental priorities, or changes in strategy. In high-stakes enterprise settings, such as compliance reporting, financial filings, and policy drafting, these contradictions are more than nuisances. They can lead to regulatory violations, operational inefficiencies, unclear ownership of responsibilities, or costly business errors.
@@
LLMs still struggle to reason over such conflicts or identify them reliably. Retrieval-augmented generation (RAG) pipelines aim to ground outputs in external evidence, but unresolved conflicts in retrieved documents like outdated policies and contradictory department guidelines often slip through unchecked. When LLMs answer enterprise queries without retrieval grounding, they hallucinate heavily, sometimes fabricating procedures or misrepresenting standards [1]. Even enterprise-domain RAG systems show high hallucination rates, with over 17% of benchmark queries producing flawed outputs due to irrelevant or conflicting evidence [2]. When contradictions in the input are not resolved, generation models tend to merge them, producing misleading or risky content.
@@
To benchmark and mitigate these failures, we need controlled generation of contradictions. Unlike natural contradictions that occur by accident, we focus on intentional contradiction injection, i.e., generating enterprise-style documents that contain subtle, structured contradictions by design. These contradictions may occur within a document (self-contradictions) or across documents (pairwise contradictions). This controllability is essential for evaluating whether models can detect, resolve, or reason over contradictions when needed. It also enables robust stress-testing of RAG pipelines and contradiction-aware generation systems.
@@
Contradiction detection remains difficult for both humans and machines. Many conflicts are subtle, context-dependent, or buried in long documents, and there is a scarcity of domain-specific datasets. Even state-of-the-art models like GPT-4 [3] and LLaMAv2 [4] perform only slightly better than chance on contradiction detection tasks [5]. These challenges are compounded by the difficulty of defining ground truth in enterprise corpora. Our framework incorporates a human-in-the-loop setup for quality assurance and for annotating contradiction pairs to support robust evaluation.
I-A Challenges in Contradiction Benchmarking
@@
    Complex Enterprise Language: Existing contradiction benchmarks often use simplified sentences, whereas real enterprise documents (policies, compliance reports, HR manuals) contain jargon-heavy, multi-clause structures.
@@
    Limited Contradiction Types: Most datasets only handle binary contradiction/entailment, overlooking enterprise-specific cases like temporal mismatches, numerical inconsistencies, conflicting responsibilities, or strategy reversals.
@@
    Lack of Cross-Document Contradictions: Prior work emphasizes single-document conflicts, while in enterprises, contradictions frequently span multiple policies, reports, or department outputs.
@@
    Manual Contradiction Detection is Time-Intensive: Identifying contradictions across large corpora of policies, contracts, or filings requires tedious line-by-line checks, making manual detection impractical at scale.
I-B Our Contribution
To address these challenges, we introduce a novel multi-agent contradiction-aware benchmark generation framework that supports generation of realistic enterprise-style documents:
@@ I-B Our Contribution
    Realistic Enterprise Language: We generate synthetic documents that mirror enterprise tone, structure, and metadata. The framework supports customizable domains (e.g., HR, compliance, finance, strategy) and ensures quality through perplexity checks, LLM scoring, and human validation.
@@
    Rich Contradiction Taxonomy: We explicitly model six types of contradictions (Table I) going beyond simple negation to capture conflicts common in enterprise workflows, including temporal, numerical, and authority-based inconsistencies.
@@
    Pairwise Contradictions: We enable controlled creation of both self-contradictions (within a single document) and pairwise contradictions (across documents), simulating real-world enterprise conflicts between departments or policies.
@@
    Human-in-the-Loop Supervision: Human oversight is embedded at multiple stages, from verifying fluency and realism to resolving ambiguous contradiction cases, ensuring that generated corpora reflect enterprise-level fidelity and human judgment complexity.
To our knowledge, this is one of the first multi-agent frameworks for generating enterprise-style documents with controllable contradictions. The resulting synthetic corpora allow researchers to evaluate how well systems detect and reason over conflicts, providing a foundation for more reliable enterprise document generation, retrieval filtering, and contradiction-aware evaluation.
@@
TABLE I: Taxonomy of Contradiction Types
II Background and Related Work
II-A Contradictions in Generation and Retrieval
@@
Natural Language Inference (NLI) classifies the logical relationship between two sentences as entailment, neutrality, or contradiction. Contradiction detection focuses on identifying when two statements cannot both be true under the same interpretation. Contradictions appear either within a single document (self-contradictions) or across multiple documents covering related content (pairwise contradictions). In retrieval-augmented generation (RAG), unresolved conflicts like outdated policies, overlapping compliance guidelines, or inconsistent departmental reports can cause models to produce factually incorrect or logically inconsistent outputs. Left unfiltered, these contradictions represent a high-risk form of hallucination in enterprise settings. Consider the following example, where two HR policies specify different waiting periods for remote work eligibility.
Example: Contradictory Policies Input Documents: Policy A: “Employees are eligible for remote work after 30 days of service.” Policy B: “Employees are eligible for remote work after 90 days of service.” Query: “When can new employees start working remotely?” LLM Output: “New employees are eligible for remote work after 30 to 90 days, depending on circumstances.”
The model merges contradictory rules instead of resolving them, producing an ambiguous and non-compliant answer. ContraGen addresses this by intentionally injecting such conflicts in synthetic enterprise documents, enabling systematic evaluation of whether models can detect, localize, and resolve them.
Contradictions in RAG are increasingly studied as both hallucinations and knowledge conflicts, tied to broader challenges of factuality and reliability in LLM outputs. Zhang and Zhang [6] review hallucination mitigation in retrieval-augmented LLMs, showing that contradictions often stem from retrieval failures (outdated or conflicting sources) and generation deficiencies (context conflict, misalignment), making them one of the most risk-prone forms of hallucination in RAG. Wang et al. [7] propose a framework for evaluating knowledge conflicts between parametric and retrieved information, finding that while LLMs can detect conflicts, they struggle to localize them or provide distinct answers. Together, these studies suggest that contradictions in enterprise and compliance documents are not incidental errors but structural bottlenecks in retrieval and generation, motivating benchmarks that intentionally embed self and pairwise contradictions.
@@
Despite growing interest, existing contradiction benchmarks are poorly suited for stress-testing generative systems. Early work in NLI introduced datasets like SNLI [8], MNLI [9], and ANLI [10], but these focus on isolated sentence pairs and lack the linguistic and structural complexity of enterprise documents. Recent datasets such as ContraDoc [5] and ECON [11] move toward document- and retrieval-scale evaluation. ContraDoc showed that even GPT-4 struggles with subtle internal inconsistencies in long documents, and trained annotators also miss many contradictions, highlighting that even human review is unreliable at scale. ECON distinguishes between factoid and answer conflicts, showing that models often default to internal priors when evidence is inconsistent. Gokul et al. [12] extend this line of work to synthetic RAG pipelines by generating controlled contradictions and testing LLMs as context validators, but they consider only three types: self, pairwise, and conditional contradictions.
These works highlight progress toward contradiction-aware evaluation but leave open the question of how to build realistic, domain-specific corpora that reflect the retrieval conflicts faced in enterprise applications—where overlapping reports, shifting KPIs, and conflicting compliance requirements are the norm.
Figure 1: ContraGen - Overall Workflow
II-B Multi-Agent and Generative Systems
Recent work highlights the growing role of multi-agent architectures in improving reliability and structure in LLM-based generation. Musumeci et al. [13] show that agent specialization—splitting tasks such as template decomposition, retrieval, and iterative drafting—can reduce hallucinations and produce more consistent public administration documents. Xie et al. [14] survey multimodal agents, emphasizing the need for perception, planning, and memory as core components of scalable deployment. Wang et al. [15] demonstrate that structured communication and hierarchical refinement in multi-agent systems can improve task performance in collaborative settings. M-RAG [16] leverages multi-agent reinforcement learning over database partitions, showing how specialization in retrieval and refinement can reduce noise and improve reliability.
Building on these insights, we design a multi-agent framework that applies specialization to contradiction-aware corpus generation. By decoupling generation, contradiction injection, and evaluation into separate roles, our approach provides finer control over the types of contradictions introduced and automates quality checks, directly addressing the gap in current benchmarks, which lack realistic, controllable corpora that reflect the retrieval conflicts common in enterprise and compliance workflows.
III System Architecture
@@
ContraGen is a modular multi-agent system designed for generating and evaluating enterprise documents enriched with controlled contradictions. It supports document creation, contradiction injection, fluency validation, hybrid contradiction mining, and retrieval-based fact-checking. All agents communicate via Pyro4-based remote procedure calls and are coordinated by a central orchestrator. The overall architecture is shown in Figure 1.
ContraGen comprises three specialized agents, each addressing a distinct stage in the document generation and evaluation pipeline. The Content Generator Agent produces structured metadata and base documents for selected subdomains, injects self- and pairwise contradictions via instruction-tuned prompts, and validates fluency using relative perplexity scores while logging detailed metadata, including conflict pairs. The Contradiction Mining Agent employs a hybrid approach that combines a pretrained Natural Language Inference (NLI) model with an LLM judge, leveraging semantic similarity filtering to identify relevant sentence pairs and detect both self and cross-document contradictions. A human-in-the-loop complements this process by verifying that injected contradictions actually manifest as contradictions in context, manually adding injected cases that the system fails to detect, and annotating detections. Since injected conflicts can interact with surrounding content and create additional, unintended inconsistencies, human reviewers also provide contextual judgments to ensure that the evaluation set faithfully represents both intended and emergent contradictions. Finally, the Retrieval Verifiability Agent evaluates whether conflicting statements can be externally verified using publicly available evidence, enabling a two-track evaluation framework that distinguishes between retrieval-verifiable and retrieval-resistant contradictions.
@@
III-A Contradiction-Aware Content Generation Agent
III-A1 Metadata-Driven Base Document Generation
To initiate generation, we construct a realistic organizational profile that serves context for generating metadata. The scope of enterprise domain is organized into six categories. Given the organization description and subdomain structure, metadata is generated containing title, topic, date, department, location, document type, and authority level. This structured metadata serves as the source for multi-paragraph enterprise documents generated in assertive, policy-oriented language (see Prompt: Base Document Generation in Appendix A-A). We discuss the exact organization profile and domain-subdomain structure using in our framework in section IV.
III-A2 Perplexity-Based Fluency Control
@@
where $P(w_{i}\mid w_{<i})$ is the probability assigned to the $i$-th token given its preceding context $w_{<i}$, and $N$ is the total number of tokens in the document. A lower perplexity indicates that the model finds the text more predictable (i.e., fluent) under its learned distribution. To ensure fluency preservation after contradiction injection, the perplexity of both the base and contradicted document is computed using pretrained language model GPT2 [17]. Let $PPL_{\text{base}}$ denote the perplexity of the base document and $PPL_{\text{contr}}$ the perplexity after contradiction injection. If the perplexity $PPL>22$, the generated document is rejected and re-tested. The relative change is defined as:
@@
III-A3 Contradiction Injection
To embed a self-contradiction (see Figure 2) , a target sentence is first selected using an LLM prompt that identifies a factual, specific, and authoritative statement. A second prompt with few-shot contradiction type examples generates a direct and confident contradiction statement, explicitly avoiding hedging language (e.g., “however,” “may,” “sometimes”). Finally, a third prompt blends the contradiction seamlessly into the document, preserving coherence, tone, and style. The prompts used are mentioned in Appendix A-B.
@@ III-A3 Contradiction Injection
We conduct a human evaluation of generated documents containing injected contradictions using a 5-point Likert scale. The results show strong average ratings for fluency (4.2), specificity (4.7), and coherence (4.5), indicating the documents were well-written, detailed, and structurally consistent. Legitimacy scored slightly lower (3.5), suggesting that some documents lacked the authority expected of authentic enterprise-style texts. Contradictions were detected in only 54% of the cases, and when identified, were rated as subtle and realistic. Although our injected contradictions are naturalistic, they remain challenging for even human reviewers to reliably detect.
III-B Contradiction Mining Agent
Long-form documents may contain subtle inconsistencies which makes contradiction identification challenging for both humans and models. The Contradiction Mining Agent identifies high-likelihood contradiction pairs, which are then validated by human annotators. This agent ensures: 1) benchmark quality by verifying non-trivial, detectable contradictions and 2) provides a baseline for assessing RAG systems’ ability to recognize conflicts before generation.
NLI models are fast and widely used but they often struggle with spurious correlations, brittleness, and the difficulty of capturing nuanced reasoning. LLMs, on the other hand, offer richer reasoning but are computationally expensive. To identify contradictions at scale across documents, we combine fast retrieval-style filtering with multi-stage entailment reasoning. Our approach avoids exhaustive $\mathcal{O}(n^{2})$ sentence pair comparison.

### v2

# The First Workshop on Generative and Protective AI for Content Creation

@@ Abstract
Retrieval-Augmented Generation (RAG) integrates large language models (LLMs) with external sources, but unresolved contradictions in retrieved evidence often lead to hallucinations and legally unsound outputs. Benchmarks currently used for contradiction detection lack domain realism, cover only limited conflict types, and rarely extend beyond single-sentence pairs, making them unsuitable for legal applications. Controlled generation of documents with embedded contradictions is therefore essential: it enables systematic stress-testing of models, ensures coverage of diverse conflict categories, and provides a reliable basis for evaluating contradiction detection and resolution. We present a multi-agent contradiction-aware benchmark framework for the legal domain that generates synthetic legal-style documents, injects six structured contradiction types, and models both self- and pairwise inconsistencies. Automated contradiction mining is combined with human-in-the-loop validation to guarantee plausibility and fidelity. This benchmark offers one of the first structured resources for contradiction-aware evaluation in legal RAG pipelines, supporting more consistent, interpretable, and trustworthy systems.
1 Introduction
@@
Large language models (LLMs) are increasingly used to draft legal documents, synthesize policies, and answer regulatory questions. These tasks differ significantly from general-purpose generation or factoid question answering. Legal text is dense, contextual, and often contains contradictions  [^1] In our case, contradictions may also correspond to policy violations or IP infringements., either introduced inadvertently, or arising from evolving regulations, overlapping jurisdictions, or competing organizational priorities. In high-stakes domains like law, compliance, and policy drafting, such contradictions are not just nuisances. They can lead to regulatory non-compliance, contractual disputes, and unclear ownership of intellectual property.
@@
LLMs still struggle to reason over these conflicts or reliably identify them. Retrieval-augmented generation (RAG) pipelines attempt to ground outputs in external evidence, but unresolved conflicts in that evidence, such as superseded laws or inconsistent interpretations, often pass unchecked into the output. LLMs answering legal queries without retrieval grounding hallucinated between 69% and 88% of the time, fabricating statutes or misapplying case law (Dahl et al., 2024). Even legal-domain RAG systems hallucinate on over 17% of benchmark queries (Magesh et al., 2025), largely due to the retrieval of irrelevant or conflicting sources. When contradictions in input evidence go unresolved, generation models often merge them, producing legally unsound and potentially risky outputs.
@@
To benchmark and mitigate these failures, we need controlled generation of contradictions. Unlike natural contradictions that emerge as side-effects or errors, we focus on intentional contradiction injection, i.e., the ability to generate legal-style documents that contain subtle, structured contradictions by design. These contradictions may occur within a document (self-contradictions) or across documents (pairwise contradictions). This kind of controllability is essential for evaluating whether models can detect, resolve, or reason over contradictions when needed. It also opens the door to robust stress-testing of RAG pipelines and contradiction-aware generation systems.
@@
Contradiction detection remains difficult for both humans and machines due to the subtle, context-dependent nature of many conflicts and the scarcity of domain-specific datasets. Even state-of-the-art models like GPT-4 (OpenAI, 2024) and LLaMAv2 (Touvron et al., 2023) perform only slightly better than chance (Li et al., 2023). These challenges are compounded by the difficulty of defining ground truth in long-form legal texts. Our framework incorporates a human-in-the-loop setup for post-generation QA and for annotating contradiction pairs to support robust evaluation.
1.1 Challenges in Contradiction Benchmarking
@@
    Unrealistic Legal Language: Current contradiction benchmarks often use overly plain or generic sentences, whereas real legal documents contain formal, complex sentence structures, and use intricate and nuanced language.
@@
    Limited Contradiction Types: Most datasets focus on binary contradiction/entailment, overlooking complex types like temporal misalignment, reversals in obligations, or conflicts in legal authority - common in legal documents.
@@
    Lack of Cross-Document Contradictions: Most prior work focuses on single-document contradictions, while real conflicts often span multiple contracts, policies, or filings.
@@
    Manual Contradiction Detection is Time-Intensive: Detecting contradictions across contracts or policies is slow and error-prone, often requiring line-by-line review.
1.2 Our Contribution
To address these challenges, we introduce a novel multi-agent contradiction-aware benchmark generation framework that supports controllable generation of legal-style documents:
@@ 1.2 Our Contribution
    Realistic Legal Language: We generate synthetic documents that mirror real legal tone, structure, and metadata. The framework supports customizable domains and subdomains, and quality is evaluated via perplexity, LLM assessment, and human validation.
@@
    Rich Contradiction Taxonomy: We explicitly model six types of contradictions (Table 1) that goes beyond simple negation and captures the kinds of conflicts legal professionals routinely confront.
@@
    Pairwise Contradictions: We enable the controlled creation of documents with both self-contradictions (contradictions present within the document) and pairwise-contradictions (contradictions present across documents) reflective of legal drafting norms.
@@
    Human-in-the-Loop Supervision: Human validation is embedded throughout the pipeline, from verifying contradiction realism and fluency, to resolving ambiguous contradiction pairs, ensuring that generated corpora maintain legal fidelity and that detection benchmarks reflect real-world judgment complexity.
To our knowledge, this is one of the first multi-agent generation framework for creating legal-style documents with controllable contradictions, designed to simulate the types of conflicts that arise in real-world legal, regulatory, and compliance workflows. These synthetic corpora allow us to evaluate how well systems detect and reason over contradictions, providing a foundation for more reliable document generation, retrieval filtering, and contradiction-aware evaluation in high-stakes domains.
@@
Table 1: Taxonomy of Contradiction Types
2 Background and Related Work
2.1 Contradictions in Generation and Retrieval
@@
Natural Language Inference (NLI) involves determining the logical relationship between two sentences as entailment, neutrality, or contradiction. Contradiction detection is a focused subtask that identifies when two statements cannot simultaneously be true under the same interpretation. Contradictions can occur either within a single document (self-contradictions) or across multiple documents on related content (pairwise contradictions). In retrieval-augmented generation (RAG), unresolved conflicts like outdated statutes, overlapping jurisdictions, or inconsistent interpretations can lead models to produce factually incorrect or logically inconsistent outputs. Left unfiltered, these contradictions represent a high-risk form of hallucination in legal and policy settings.
@@
Despite growing interest, existing contradiction benchmarks are poorly suited for stress-testing generative systems. Early work in NLI introduced datasets like SNLI (Bowman et al., 2015), MNLI (Williams et al., 2017), and ANLI (Williams et al., 2020), but these focus on isolated sentence pairs and lack the linguistic and structural complexity of real documents. More domain-specific resources such as ContractNLI (Koreeda and Manning, 2021) are limited to clause-level reasoning. ContraDoc (Li et al., 2023) and ECON (Jiayang et al., 2024) move toward document- and retrieval-scale evaluation. ContraDoc showed that even GPT-4 struggles with subtle internal inconsistencies in long documents, and that trained annotators also miss many contradictions. ECON distinguishes factoid and answer conflicts, showing that models often default to internal priors when evidence is inconsistent. Gokul et al. (2025) extend this work to synthetic RAG pipelines by generating controlled contradictions and testing LLMs as context validators, but consider only three types: self, pairwise, and conditional.
Figure 1: LegalWiz - Workflow
2.2 Multi-Agent and Generative Systems
Recent literature has explored content generation through multi-agent architectures. Musumeci et al. (2024) propose a multi-agent architecture that automates the drafting of semi-structured public administration documents by decomposing templates, retrieving relevant information, and generating text iteratively to reduce hallucinations, demonstrating how agent specialization can yield more structured and reliable outputs. Xie et al. (2024) survey large multimodal agents and emphasize perception, planning, and memory as essential components for real-world deployment. Wang et al. (2025) introduce TalkHier, a collaborative LLM-MA framework that improves task performance through structured communication and hierarchical refinement.
Current benchmarks leave open a critical gap in building realistic, controllable corpora that simulate the retrieval conflicts common in legal and compliance settings. Our agentic framework allows us to decouple generation, contradiction injection, and evaluation into specialized roles, making it easier to control the types of contradictions introduced and automate quality checks for the creation of benchmark-ready corpora.
3 System Architecture
@@
LegalWiz is a modular multi-agent system designed for generating and evaluating legal documents enriched with controlled contradictions. It supports document creation, contradiction injection, fluency validation, hybrid contradiction mining, and retrieval-based fact-checking. All agents communicate via Pyro4-based remote procedure calls and are coordinated by a central orchestrator. The overall architecture is shown in Figure 1.
LegalWiz integrates three coordinated agents spanning document generation to verification. The Content Generator Agent creates structured metadata and base legal documents, injects self- and pairwise contradictions via instruction-tuned prompts, and validates fluency using relative perplexity while logging conflict metadata. The Contradiction Mining Agent combines an NLI model with an LLM judge, using semantic similarity filtering to detect intra- and inter-document conflicts. A human-in-the-loop review refines low-confidence cases. The Retrieval Verifiability Agent then assesses whether conflicting statements are externally verifiable, distinguishing retrieval-verifiable from retrieval-resistant contradictions.
@@
3.1 Contradiction-Aware Content Generation Agent
3.1.1 Metadata-Driven Base Document Generation
To initiate generation, we construct a realistic organizational profile that serves context for generating metadata. The scope of legal domain is organized into five categories. Given the organization description and subdomain structure, metadata is generated containing title, topic, date, department, location, document type, and authority level. This structured metadata serves as the source for multi-paragraph legal documents generated in assertive, policy-oriented language (see Prompt: Base Document Generation in A.1). We discuss the exact organization profile and domain-subdomain structure using in our framework in section 4.
3.1.2 Perplexity-Based Fluency Control
@@
where $P(w_{i}\mid w_{<i})$ is the probability assigned to the $i$-th token given its preceding context $w_{<i}$, and $N$ is the total number of tokens in the document. A lower perplexity indicates that the model finds the text more predictable (i.e., fluent) under its learned distribution. To ensure fluency preservation after contradiction injection, the perplexity of both the base and contradicted document is computed using pretrained language model GPT2 (Radford et al., 2019). Let $PPL_{\text{base}}$ denote the perplexity of the base document and $PPL_{\text{contr}}$ the perplexity after contradiction injection. If the perplexity $PPL>22$, the generated document is rejected and re-tested. The relative change is defined as:
@@
3.1.3 Contradiction Injection
To embed a self-contradiction (see Figure 2) , a target sentence is first selected using an LLM prompt that identifies a factual, specific, and authoritative statement. A second few-shot prompt generates a direct and confident contradiction statement, explicitly avoiding hedging language (e.g., “however,” “may,” “sometimes”). Finally, a third prompt blends the contradiction seamlessly into the document, preserving coherence, tone, and style. The prompts used are mentioned in the appendix A.2).
@@ 3.1.3 Contradiction Injection
We conduct a human evaluation of generated documents containing injected contradictions using a 5-point Likert scale. The results show strong average ratings for fluency (4.2), specificity (4.7), and coherence (4.5), indicating the documents were well-written, detailed, and structurally consistent. Legitimacy scored slightly lower (3.5), suggesting that some documents lacked the authority expected of authentic legal texts. Contradictions were detected in only 43% of the cases, and when identified, were rated as subtle and realistic. Although our injected contradictions are naturalistic, they remain challenging for even human reviewers to reliably detect.
3.2 Contradiction Mining Agent
Long-form legal documents often contain subtle inconsistencies which makes contradiction identification challenging for both humans and models. The Contradiction Mining Agent identifies high-likelihood contradiction pairs, which are then validated by human annotators. This agent ensures: 1) benchmark quality by verifying non-trivial, detectable contradictions and 2) provides a baseline for assessing RAG systems’ ability to recognize conflicts before generation.
NLI models are fast and widely used but they often struggle with spurious correlations, brittleness, and the difficulty of capturing nuanced reasoning. LLMs, on the other hand, offer richer reasoning but are computationally expensive. To identify contradictions at scale across synthetic legal documents, we combine fast retrieval-style filtering with multi-stage entailment reasoning. Our approach avoids exhaustive $\mathcal{O}(n^{2})$ sentence pair comparison.